OpenStudy (anonymous):

solve please..(2x2 – 23x + 63) ÷ (x – 7)

OpenStudy (diyadiya):

Factorise \[2x^2 – 23x + 63=(x-7)(2x-9)\] \[\frac{(x-7)(2x-9)}{x-7}\] \[=(2x-9)\]